If y varies jointly with x cubed and the square root of z, find the constant of variation if y=12 when x=4 and z=49
Thanks in advance for all of your help

Answer by stanbon(75887)

If y varies jointly with x cubed and the square root of z,
------------------
y = kx^3*sqrt(z)
---------
find the constant of variation if y=12 when x=4 and z=49
Solve for "k":
12 = k*4^3*sqrt(49)
----
12 = k*64*7
----
k = 0.0268
==============
Cheers,
Stan H.
==============